# Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions

### v2

@@ II.2.4 Hadronization via coalescence
where $\mu_{12}$ is the reduced mass of the two quarks undergoing coalescence and $T$ is the temperature of the medium. The numerator is associated with the probability of forming a meson in $r$-space, which of course increases with the width $\sigma_{r}$. On the other hand, this increase is compensated by the probability in momentum space where the effective width is given by $\sigma_{p}=1/\sigma_{r}$, hence decreasing with increasing $\sigma_{r}$. From the denominator of Eq. 8 we can see that there is another scale that plays a role and is given by $1/2\mu T$. This represents the width of the distribution of relative momentum of the coalescing objects. In the infinite temperature limit that would give a distribution in momentum space asymptotically flat, and the yield would become independent on the wave function width. Increasing $\sigma_{r}$ one increases the probability of coalescence in r-space but this is exactly compensated by the shrinking of the Wigner hadron function in momentum space. Apart from this extreme ideal case, the increase of the width $\sigma_{r}$ leads to an increase that quantitatively depends on the comparison between the width of the hadron wave function $\sigma_{r}$ and the width of the thermal distribution $1/2\mu T$. For $\sigma_{r}^{2}<<(2\mu T)^{-1}$ the yield is strongly dependent on the size of the hadron wave function, while for $\sigma_{r}^{2}>>(2\mu T)^{-1}$ the yield is nearly independent on the hadron wave function. For the specific case of charm meson and baryons, $\sigma_{r}^{2}(\mathrm{D^{0}})=0.65\,\mathrm{f}m^{2}$ and $(2\mu T)^{-1}=0.5\mathrm{\,}fm^{2}$, hence an increase of $<r^{2}>$ by a factor of two implies an increase of the yield of about 30%. Smaller baryons like $\Omega_{c}$ have the relative widths $\sigma_{r1}^{2}(\Omega_{c})=0.34\,\mathrm{f}m^{2}$ and $\sigma_{r2}^{2}(\Omega_{c})=0.14\,\mathrm{f}m^{2}$ and $(2\mu_{ss}T)^{-1}=0.6\mathrm{\,}fm^{2}$ and $(2\mu_{(ss)c}T)^{-1}=0.6\mathrm{\,}fm^{2}$ such that an increase of $<r^{2}>_{\Omega_{\mathrm{c}}^{0}}$ by a factor of 2 induces an increase of the yield of about a factor 2.2. Therefore, hadrons that have a large mean square radius and large mass are in general not very sensitive to the parameter $\sigma_{r}$, if the relation $\sigma_{p}=1/\sigma_{r}$ is employed. However, on top of this standard coalescence feature, one has to consider the constraint that $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ strongly damp such dependence for the D and $\Lambda_{\mathrm{c}}^{+}$ hadrons, which dominate the overall charm-hadron yield [74]. The Catania coalescence plus fragmentation model was first applied to AA collisions prediction of $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ both at RHIC and LHC energy [37], but soon after it led to a quite good prediction of the $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ yield ratio in pp collisions, as shown in the right panel of Fig. 6. It also provided within the current uncertainties also a reasonable predictions for the $\Xi_{\mathrm{c}}^{0}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$ production, see also Fig.s 15 and 16 in Sec.III.C.
@@
Other coalescence approaches: In other approaches [34, 88], applied to AA collisions and based on a formulation formally very similar to Eq. 6 and 7, the constraint to have $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ is also implemented, but it is achieved artificially increasing the width $\sigma_{ri}$ of hadrons by a constant factor, resulting in radii about a factor 1.8 larger than the one employed by the Catania model [37] that relies on the quark model prediction in Ref. [86, 87]. This choice should also affect the width in momentum space modifying the $p_{\mathrm{T}}$ dependence of the spectra. A study of such an effect, and a direct comparison to the Catania model, would be instructive but has not yet been performed. A refinement of the coalescence approach is to consider the different wave function of $l=1$ states, which leads to a non-Gaussian Wigner function [88], different from that of the Catania model. This is done also in the hadronization mechanism implemented in the Parton Hadron String Dynamics (PHSD) [89].
@@
The PHSD model provides a realistic simulation of the entire heavy-ion collision and it also implements a coalescence approach to hadronization of heavy quarks [89]. The quark masses are distributed according to the spectral function of a dynamical quasi-particle model (DQPM) [90], very similar to the QPM used in Catania to describe the heavy-quark dynamics. Then, as in the Catania model, the probability of coalescence is evaluated by a superposition of the quark distribution with the Wigner hadron wave function, but considering both S- and P-wave states. Differently from the Catania model, it does not implement the constraint $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ and the coalescence process is not instantaneous but occurs during the fireball expansion when the local energy density spans the interval $0.3<\epsilon<0.9\,\mathrm{G}eV/fm^{3}$. Hadronization for heavy flavor within the PHSD approach has been extensively applied to study charm dynamics in AA collisions; only very recently it has been done an extension to pp collision studying the bottomonium production [91].
@@
Finally, a more simplified approach to hadronization by coalescence is the Quark re-Combination Model (QCM) which, though based on the idea of quark recombination, has several features that are different from the other approaches discussed above. The model is developed in one dimension representing the transverse momentum. The coordinate space is not considered. The model does not employ as light-quark sources quark distribution functions as thermal+flow distributions and/or resulting from a dynamical study of heavy quark dynamics. Rather, it extracts them from a fit to $\pi$, K, and D mesons [92, 35]. Furthermore, hadronization is occurring by coalescence at all momenta, without any contribution from fragmentation. Once the u, d, s, and c quark distribution functions are extracted from the fit, they are used as a source to compute the spectra as a function $p_{\mathrm{T}}$ of all hadron species: p, $\Lambda$, $\phi$, $\Omega$, $\Lambda_{\mathrm{c}}^{+}$ etc… The role of Wigner hadron function in the standard recombination approach is substituted by a recombination function that is given by a delta function, which imposes quarks with the same velocity to recombine. The ratio between vector and pseudoscalar meson yields is fixed by a parameter $R_{V/P}=3/2$ indicated as thermal weight because it corresponds to the value of $\mathrm{D}^{*}/D\simeq 3\,(M_{D}/M_{D^{*}})^{3/2}exp[-(M_{D}-M_{D^{*}})/{\it T_{H}}]\simeq 3/2={\it R_{V/P}}$, if one assumes a thermal production at a temperature $\mathrm{\it T_{H}}=170-175\,MeV$, as done also in [93] in the context D meson production in $\pi N$ collisions, see Sect. III.B. This is the same assumption exploited in the Catania coalescence model even if in this case the temperature, $\mathrm{\it T_{H}}=165\,MeV$ is the one of the hadronizing medium that determines the light and strange quark thermal distribution, while in QCM the concept of temperature does not really play a direct role because the quark distribution functions are deduced from a fit to $\pi$, K, and D mesons experimental data. A similar ratio parameter is used to fix the ratio between the baryons of the octet ($J^{P}=(1/2)^{+}$) and decuplet ($J^{P}=(3/2)^{+}$) by the parameter $R_{O/D}=2$.
@@
Resonance recombination model: Another main approach to charm hadron formation is the Resonance Recombination Model (RRM) of TAMU group where hadronization is described by scattering of thermal light quarks with heavy quarks into broad D-meson like resonances [63]. This model is also based on the idea that hadronization occurs by a recombination of quarks in the medium, but the process is modelled like a scattering of two particles into a hadron resonance with a finite width $\Gamma$. The meson spectrum is written in terms of the quark and anti-quark distribution as:
$\frac{dN_{M}}{d^{2}p_{\mathrm{T}}dy}=\int\frac{d^{3}\vec{p}_{1}d^{3}\vec{p}_{2}}{(2\pi^{3})}f_{q}(\vec{x},\vec{p}_{1})f_{\overline{q}}(\vec{x},\vec{p}_{2})\frac{\sigma_{M}(s)}{M\Gamma}v_{rel}(\vec{p}_{1},\vec{p}_{2})\delta^{3}(\vec{p}-\vec{p}_{1}-\vec{p}_{2}),$ (9)
@@
where $\sigma(s)$ is the relativistic resonance cross section (Breit-Wigner). We can envisage from Eq. 9 and Eq. 6 that RRM approach shares with the coalescence approach several features because the main underlying process consists in adding momenta of particles coming from the medium and the relative momenta of quarks determine the formation probability according to the values of $\sigma_{M}(s)\,v_{rel}$. On the other hand, the formation of the hadron is not directly related to the wave function width. However, the main features of modifying $R_{\mathrm{A}A}(p_{\mathrm{T}})$ from charm to D meson inducing an enhancement of D-meson $R_{\mathrm{A}A}$ at $p_{\mathrm{T}}\gtrsim m_{\mathrm{Q}}$ accompanied by an enhancement of the elliptic flow are similar to those resulting from the coalescence approach. The main difference with respect to coalescence is in the momentum dependence of the probability of hadron formation by recombination. Also in RRM there are two competing hadronization processes, with independent fragmentation applied to the charm quark that does not recombine according to Eq. 9. However, the recombination in RRM is dominant over fragmentation in a wider $p_{\mathrm{T}}$ range with respect to coalescence [94]. The RRM approach has been more recently applied also to the $\Lambda_{\mathrm{c}}^{+}$ production, which is modelled as a two-step process: the formation of a diquark according to Eq. 9 followed by a scattering process between the diquark and the charm quark. An important ingredient introduced in recent years in the modelling of coalescence in the TAMU model are space-momentum correlations between the recombining quarks. Given that recombination is more probable for quarks that are close in both momentum and space, the correlation between the heavy-quark momentum vector and their spatial position in the fireball, induced by the interplay of the heavy-quark interaction in the medium and the medium expansion, influences significantly the heavy-flavour hadron $R_{\mathrm{AA}}$, $v_{\mathrm{2}}$, and the relative particle-species abundances. Furthermore, being the hadronization modelled like a scattering process $2\rightarrow 1$ going into an off-shell hadron it is possible to conserve the 4-momentum and it has been shown that under the assumption of local thermal equilibrium the D meson spectrum converges to the equilibrium thermal spectrum. A more thorough comparison to the coalescence approach has indeed to consider that qualitatively coalescence is quite similar to RRM because the pair going into coalescence to form a D meson have an invariant-mass distribution around the D-meson mass with a finite width, so that energy is conserved with respect to an off-shell D-meson-like resonance also in a coalescence approach. The difference is that however in RRM the exact thermal equilibrium is constrained while in coalescence one has simply an invariant-mass distribution regulated by the relative quark momentum distribution in the medium and its interplay with the wave function width. This aspect has never been pointed out: a direct comparison of coalescence, RRM, and the diquark-quark in medium hadronization, recently developed by the Turin group within the Powlang model discussed below, could be interesting.
@@
where the mass $M$ of quark and di-quarks are taken in the range of the constituent quark picture: $m_{\mathrm{u}/d}=0.33$ GeV, $m_{\mathrm{s}}=0.5$ GeV, $m_{(\mathrm{u}d)_{0}}=0.579$ GeV, $m_{(\mathrm{s}l)_{0}}=0.804$ GeV, $m_{(\mathrm{s}s)_{1}}=1.0936$ GeV; while $g_{s}$ is the spin degeneracy and $g_{I}$ the isopin degeneracy. The recombination occurs if the invariant mass of the pair, $M_{\cal C}$, is larger than the mass of the lightest charmed hadron. Then the cluster decays into a charmed hadron that has the same baryon number and strangeness of the cluster or, for large invariant masses, $M_{\cal C}>M_{max}\simeq 3.8\,\mathrm{G}eV$, it hadronize via string-fragmentation according to PYTHIA 6.4. Therefore, differently from the coalescence approach, the probability of hadronizing via recombination is not determined by a hadron wave function with a finite width in momentum space (typically of the order of a few hundred MeV), but rather by the formation-probability of clusters of quite large invariant mass undergoing to decay. Instead, in a coalescence approach the invariant-mass distribution from the recombination of a light (strange) quark and a heavy quark can be quite narrower due to the constraint of the wave function. Another difference is the assumption of a thermal population of di-quarks that in the modelling acquire also a hydrodynamical radial and elliptic flows. The assumption in the model of local colour neutralization leads to a strong space-momentum correlation, which provides a substantial enhancement of the elliptic flow of the final-state charmed hadrons, affecting both their momentum and their angular distributions. This hadronization scheme has been applied to AA collisions [96] and very recently to pp collisions [97] showing a behavior of $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ similar to the coalescence approach, apart from the tendecy to over estimate the $D_{s}^{+}$ production especially at low $p_{T}$.

### v3

@@ II.2.4 Hadronization via coalescence
where $\mu_{12}$ is the reduced mass of the two quarks undergoing coalescence and $T$ is the temperature of the medium. The numerator is associated with the probability of forming a meson in $r$-space, which of course increases with the width $\sigma_{r}$. From the denominator of Eq. 8 we can see that there is another scale that plays a role and is given by $2\mu T$. This represents the width of the distribution of relative momentum of the coalescing objects. The increase of the width $\sigma_{r}$ leads to an increase of the yield that quantitatively depends on the comparison between the width of the hadron wave function $\sigma_{r}$ and the inverse of the width of the thermal distribution $1/2\mu T$. For $\sigma_{r}^{2}<<(2\mu T)^{-1}$ the yield is strongly dependent on the size of the hadron wave function, while for $\sigma_{r}^{2}>>(2\mu T)^{-1}$ the yield is nearly independent on the hadron wave function. For the specific case of charm meson and baryons, $\sigma_{r}^{2}(\mathrm{D^{0}})=0.65\,\mathrm{f}m^{2}$ and $(2\mu T)^{-1}=0.5\mathrm{\,}fm^{2}$, hence an increase of $<r^{2}>$ by a factor of two implies an increase of the yield of about 30%. Smaller baryons like $\Omega_{c}$ have the relative widths $\sigma_{r1}^{2}(\Omega_{c})=0.34\,\mathrm{f}m^{2}$ and $\sigma_{r2}^{2}(\Omega_{c})=0.14\,\mathrm{f}m^{2}$ and $(2\mu_{ss}T)^{-1}=0.6\mathrm{\,}fm^{2}$ and $(2\mu_{(ss)c}T)^{-1}=0.6\mathrm{\,}fm^{2}$ such that an increase of $<r^{2}>_{\Omega_{\mathrm{c}}^{0}}$ by a factor of 2 induces an increase of the yield of about a factor 2.2. Therefore, hadrons that have a large mean square radius and large mass are in general not very sensitive to the parameter $\sigma_{r}$, if the relation $\sigma_{p}=1/\sigma_{r}$ is employed. However, on top of this standard coalescence feature, one has to consider the constraint that $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ strongly damp such dependence for the D and $\Lambda_{\mathrm{c}}^{+}$ hadrons, which dominate the overall charm-hadron yield [91]. This is due to the role played by the C factor described above, while in other coalescence approaches where this factor is not present the effect of modifying the width of the Wigner function has a strong impact on the baryon-to-meson yield ratio, we discuss this point just below in the dedicated paragraph. The Catania coalescence plus fragmentation model was first applied to AA collisions prediction of $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ both at RHIC and LHC energy [55], but soon after it led to a quite good prediction of the $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ yield ratio in pp collisions, as shown in the right panel of Fig. 6. It also provided within the current uncertainties also a reasonable predictions for the $\Xi_{\mathrm{c}}^{0}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$ production, see also Fig.s 15 and 16 in Sec.III.C.
@@
Other coalescence approaches: In other approaches [52, 103], applied to AA collisions and based on a formulation formally very similar to Eq. 6 and 7, the constraint to have $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ is also implemented, but it is achieved artificially increasing the width $\sigma_{ri}$ of hadrons by a constant factor, resulting in meson radii about a factor 1.8 larger than the one employed by the Catania model [55] that relies on the quark model prediction in Ref. [101, 102]. The increase of the yield due to an increase of the width $\sigma_{ri}$ is squared for baryons because the relative coordinates of the quarks enter twice for baryons. Therefore, the increase of $\sigma_{ri}$ induces a strong increase of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D}$ ratio. In particular, it has been seen that modifying $\sigma_{ri}$ to satisfy the constraint $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ results in a $\Lambda_{\mathrm{c}}^{+}/\mathrm{D}$ close to the experimental measurements. This is on the one hand similar to the Catania model, but on the other hand the effect of modifying the width of the Wigner function is very different from the Catania model where $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ is enforced independently on the $\sigma_{ri}$, and the yields of D and $\Lambda_{\mathrm{c}}^{+}$ are marginally affected by a change of the widths by some common factor, because, anyway, due to the C-factor, all low momentum charm hadronize by coalescence, as discussed at the end of the previous paragraph. We also notice that increasing $\sigma_{ri}$ as in Refs. [52, 103] also affects the width in momentum space and this should modify the $p_{\mathrm{T}}$ dependence of the spectra of baryons and mesons. A study of such an effect, and a direct comparison to the Catania model, would be instructive but has not yet been performed. Ref. [103] presents a refinement of the coalescence approach considering the different wave function for $l=1$ states, which leads to a non-Gaussian Wigner function [103], differently from that of the Catania model that currently has the same Gaussina wave function for all the states. A similar refinement is done also in the hadronization mechanism implemented in the Parton Hadron String Dynamics (PHSD) [104].
@@
The PHSD model provides a realistic simulation of the entire heavy-ion collision and it also implements a coalescence approach to hadronization of heavy quarks [104]. The quark masses are distributed according to the spectral function of a dynamical quasi-particle model (DQPM) [105], very similar to the QPM used in Catania to describe the heavy-quark dynamics. Then, as in the Catania model, the probability of coalescence is evaluated by a superposition of the quark distribution with the Wigner hadron wave function, but considering both S- and P-wave states. Differently from the Catania model, it does not implement the constraint $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ and the coalescence process is not instantaneous but occurs during the fireball expansion when the local energy density spans the interval $0.3<\epsilon<0.9\,\mathrm{G}eV/fm^{3}$. Hadronization for heavy flavour within the PHSD approach has been extensively applied to study charm dynamics in AA collisions; only very recently an extension has been done to pp collision studying the bottomonium production [106].
@@
Finally, a more simplified approach to hadronization by coalescence is the Quark re-Combination Model (QCM) which, though based on the idea of quark recombination, has several features that are different from the other approaches discussed above. The model is developed in one dimension representing the transverse momentum. The coordinate space is not considered. The model obtains the light- and heavy-quark distributions through a fit to experimental data of $\pi$, K, and D mesons [107, 53]. Furthermore, hadronization is occurring by coalescence at all momenta, without any contribution from fragmentation. Once the u, d, s, and c quark distribution functions are extracted from the fit, they are used as a source to compute the spectra as a function $p_{\mathrm{T}}$ of all hadron species: p, $\Lambda$, $\phi$, $\Omega$, $\Lambda_{\mathrm{c}}^{+}$ etc… The role of Wigner hadron function in the standard recombination approach is substituted by a recombination function that is given by a delta function, which imposes quarks with the same velocity to recombine. The ratio between vector and pseudoscalar meson yields is fixed by a parameter $R_{V/P}=3/2$ indicated as thermal weight because it corresponds to the value of $\mathrm{D}^{*}/D\simeq 3\,(M_{D}/M_{D^{*}})^{3/2}exp[-(M_{D}-M_{D^{*}})/{\it T_{H}}]\simeq 3/2={\it R_{V/P}}$, if one assumes a thermal production at a temperature $\mathrm{\it T_{H}}=170-175\,MeV$, as done also in [108] in the context D meson production in $\pi N$ collisions, see Sect. III.B. This is the same assumption exploited in the Catania coalescence model even if in this case the temperature, $\mathrm{\it T_{H}}=165\,MeV$ is the one of the hadronizing medium that determines the light and strange quark thermal distribution, while in QCM the concept of temperature does not really play a direct role because the quark distribution functions are deduced from a fit to $\pi$, K, and D mesons experimental data. A similar ratio parameter is used to fix the ratio between the baryons of the octet ($J^{P}=(1/2)^{+}$) and decuplet ($J^{P}=(3/2)^{+}$) by the parameter $R_{O/D}=2$.
@@
Resonance recombination model: Another main approach to charm hadron formation is the Resonance Recombination Model (RRM) of TAMU group where hadronization is described by scattering of thermal light quarks with heavy quarks into broad D-meson like resonances [82]. This model is also based on the idea that hadronization occurs by a recombination of quarks in the medium, but the process is modelled like a scattering of two particles into a hadron resonance with a finite width $\Gamma$. The meson phase space distribution is written in terms of the quark and anti-quark distribution as:
$\frac{dN_{M}}{d^{3}p\,d^{3}x}=\int\frac{d^{3}\vec{p}_{1}d^{3}\vec{p}_{2}}{(2\pi^{3})}f_{q}(\vec{x},\vec{p}_{1})f_{\overline{q}}(\vec{x},\vec{p}_{2})\frac{\gamma_{M}}{\Gamma_{M}}\sigma_{M}(s)v_{rel}(\vec{p}_{1},\vec{p}_{2})\delta^{3}(\vec{p}-\vec{p}_{1}-\vec{p}_{2}),$ (9)
@@
where $\sigma(s)$ is the relativistic resonance cross section (Breit-Wigner), $\gamma_{M}=E_{M}/M$ and $\Gamma_{M}$ is ans assumed width for the formed D meson which is about 0.1 GeV for meson and 0.3 GeV for baryons. We can envisage from Eq. 9 and Eq. 6 that RRM approach shares with the coalescence approach several features because the main underlying process consists in adding momenta of particles coming from the medium and the relative momenta of quarks determine the formation probability according to the values of $\sigma_{M}(s)\,v_{rel}$. On the other hand, the formation of the hadron is not directly related to the wave function width. However, the main features of modifying $R_{\mathrm{A}A}(p_{\mathrm{T}})$ from charm to D meson inducing an enhancement of D-meson $R_{\mathrm{A}A}$ at $p_{\mathrm{T}}\gtrsim m_{\mathrm{Q}}$ accompanied by an enhancement of the elliptic flow are similar to those resulting from the coalescence approach. The main difference with respect to coalescence is in the momentum dependence of the probability of hadron formation by recombination. Also in RRM there are two competing hadronization processes, with independent fragmentation applied to the charm quark that does not recombine according to Eq. 9. However, the recombination in RRM is dominant over fragmentation in a wider $p_{\mathrm{T}}$ range with respect to coalescence [109]. The RRM approach has been more recently applied also to the $\Lambda_{\mathrm{c}}^{+}$ production, which is modelled as a two-step process: the formation of a diquark according to Eq. 9 followed by a scattering process between the diquark and the charm quark. An important ingredient introduced in recent years in the modelling of coalescence in the TAMU model are space-momentum correlations between the recombining quarks. Given that recombination is more probable for quarks that are close in both momentum and space, the correlation between the heavy-quark momentum vector and their spatial position in the fireball, induced by the interplay of the heavy-quark interaction in the medium and the medium expansion, influences significantly the heavy-flavour hadron $R_{\mathrm{AA}}$, $v_{\mathrm{2}}$, and the relative particle-species abundances. Furthermore, being the hadronization modelled like a scattering process $2\rightarrow 1$ going into an off-shell hadron it is possible to conserve the 4-momentum and it has been shown that under the assumption of local thermal equilibrium the D meson spectrum converges to the equilibrium thermal spectrum.
We note that when going through coalescence objects are formed that have a total energy in the rest frame (an invariant mass) that is distributed within a finite width that is comparable to the object assumed to be formed in RRM. Hence, globally, one can say that the energy is conserved looking at the D mesons obtained as to an off-shell distribution of D-meson-like resonances with width of about 0.1-2 GeV. Furthermore, a difference remains w.r.t. RRM, where the exact thermal equilibrium is constrained, while in coalescence one has simply an invariant-mass distribution regulated by the relative quark momentum distribution in the medium and its interplay with the wave function width. This aspect has never been pointed out: a direct comparison of coalescence, RRM, and the diquark-quark in medium hadronization, recently developed by the Turin group within the Powlang model, discussed below, could be interesting.
@@
where the mass $M$ of quark and di-quarks are taken in the range of the constituent quark picture: $m_{\mathrm{u}/d}=0.33$ GeV, $m_{\mathrm{s}}=0.5$ GeV, $m_{(\mathrm{u}d)_{0}}=0.579$ GeV, $m_{(\mathrm{s}l)_{0}}=0.804$ GeV, $m_{(\mathrm{s}s)_{1}}=1.0936$ GeV; while $g_{s}$ is the spin degeneracy and $g_{I}$ the isopin degeneracy. The recombination occurs if the invariant mass of the pair, $M_{\cal C}$, is larger than the mass of the lightest charmed hadron. Then the cluster decays into a charmed hadron that has the same baryon number and strangeness of the cluster or, for large invariant masses, $M_{\cal C}>M_{max}\simeq 3.8\,\mathrm{G}eV$, it hadronize via string-fragmentation according to PYTHIA 6.4. Therefore, differently from the coalescence approach, the probability of hadronizing via recombination is not determined by a hadron wave function with a finite width in momentum space (typically of the order of a few hundred MeV), but rather by the formation-probability of clusters of quite large invariant mass subject to decay. Instead, in a coalescence approach the invariant-mass distribution from the recombination of a light (strange) quark and a heavy quark can be quite narrower due to the constraint of the wave function. Another difference is the assumption of a thermal population of di-quarks that in the modelling acquire also a hydrodynamical radial and elliptic flows. The assumption in the model of local colour neutralization leads to a strong space-momentum correlation, which provides a substantial enhancement of the elliptic flow of the final-state charmed hadrons, affecting both their momentum and their angular distributions. This hadronization scheme has been applied to AA collisions [111] and very recently to pp collisions [112] showing a behavior of $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ similar to the coalescence approach, apart from the tendecy to over estimate the $\mathrm{D_{s}^{+}}$ production especially at low $p_{T}$.